Glover Company makes three products in a single facility. These products have the following unit product costs:

Product
A B C
Direct materials $ 33.20 $ 49.70 $ 56.10
Direct labor 20.60 23.20 14.00
Variable manufacturing overhead 1.60 1.00 0.50
Fixed manufacturing overhead 13.10 8.70 9.30
Unit product cost $ 68.50 $ 82.60 $ 79.90

Additional data concerning these products are listed below.

Product
A B C
Mixing minutes per unit 1.60 0.80 0.10
Selling price per unit $ 62.00 $ 84.40 $ 77.90
Variable selling cost per unit $ 1.00 $ 1.50 $ 1.90
Monthly demand in units 2,400 3,700 1,700

The mixing machines are potentially the constraint in the production facility. A total of 6,870 minutes are available per month on these machines.

Direct labor is a variable cost in this company.

Required:

a. How many minutes of mixing machine time would be required to satisfy demand for all three products?

b. How much of each product should be produced to maximize net operating income?

c. Up to how much should the company be willing to pay for one additional hour of mixing machine time if the company has made the best use of the existing mixing machine capacity?
Answer #1
a.Total Mixing machine time required = 2400*1.60 + 3700*0.80+1700*0.10
6970 Minutes
The products should be produced on the basis of contribution margin per minute of mixing machine
A B C
Selling price 62.00 84.40 77.90
Variable costs
Direct material 33.20 49.70 56.10
Direct Labor 20.60 23.20 14.00
Variable manufacturing overhead 1.60 1.00 0.50
Variable selling cost 1.00 1.50 1.90
Contribution Margin per Unit 5.60 9.00 5.40
Milling machine minutes per unit 1.60 0.80 0.10
Contribution Margin per Minute 3.50 11.25 54.00
Production Schedule
Product Units Minutes per unit Total Minutes
A 2337.50 1.60 3740.00
B 3700.00 0.80 2960.00
C 1700.00 0.10 170.00
Total 6870.00
c.Amount per additional hour = Contribution margin = 3.5*60 = $210
As additional hour will be used to produce A